OpenStudy (anonymous):

Use the distributive property to simplify: 25(2x + 1/5)?

OpenStudy (anonymous):

50x+5

OpenStudy (anonymous):

50x+5

OpenStudy (anonymous):

can you explain how you got the answer please?

OpenStudy (anonymous):

You distribute the 25 to all the terms inside the parens: 25(2x + 1/5) = 25(2x) + 25(1/5) = 50x + 5

OpenStudy (anonymous):

According to this property, 17z + z = 17z + 1z